8. Use the information in the table below to answer the following questions Unit labor requirement U.S. Argentina Wheat 50 Beef 200400 200 (a) Does either country have an absolute advantage in the production of wheat or beef? Explain. (b) What is the opportunity cost of wheat in each country? (c) What is the opportunity cost of beef in each country(d) Analyze the comparative advantage and opportunities for trade between the US and Argentina.
Answer #1

a) US have an absolute advantage in the production of both the goods because they can produce the same good by using less amount of labor. they will be producing more, therefore absolute advantage.

b) The opportunity cost of wheat in the US is 200 / 50 = 4 labor units in Beef and in Argentina it is 400 / 200 = 2 labor units in beef.

c) The opportunity cost of Beef in the US is 50/200 = 1/4 labor in wheat and in Argentina it is 200 /400 = 1/2 labor in wheat.

d) After analysing the opportunity cost the US can produce more beef at a lower labor input that is just 1/4 and the Argentina can produce more wheat at a lower labor input that is 2. SO, the US will export beef and import wheat and Argentina will export wheat and import beef.